OpenStudy (anonymous):

Solve the system x = y/5 + 3/5 and y = 8x + 4. The x-value of the solution is

OpenStudy (anonymous):

-7/4

OpenStudy (anonymous):

??

OpenStudy (anonymous):

x=-7/4 y=-44/3

OpenStudy (anonymous):

-7/4 isnt one of my options. i have -7/3. but still im not sure how to do it.

OpenStudy (anonymous):

oops...i meant -7/3 x=(8x+4)/5 +3/5

OpenStudy (anonymous):

could you explain how you got -7/3 though? In a easy to understand way.

OpenStudy (anonymous):

x=(8x+4)/5+3/5 x=(8x+7)/5 5x=8x+7 -3x=7 x=-7/3

OpenStudy (anonymous):

thank you!